# Active learning for regression in engineering populations: A risk-informed approach

### v1

@@ 3.3 Decision Process and Active Learning
A common quality-control criterion for the machining of engineering components is maintaining a surface roughness below some threshold level $S_{\text{crit}}$, beyond which the component is no longer fit for purpose. A high-quality surface finish can significantly improve the fatigue strength, corrosion resistance and creep life of machined parts Sharkawy et al., (2014). If the surface finish is damaged via inadequate modelling or control of the machining process, the part may need to be discarded or re-machined; this has an associated utility, $C_{\text{workpiece}}$. Additionally, during the machining process an inspection can be conducted to gain access to a noisy observation of the surface roughness, with a utility $C_{\text{inspection}}$. Throughout the machining process the tool can be replaced, with a cost $C_{\text{tool}}$. Here, there is a decision, inaction, to allow the tool to continue machining with a risk to damage the workpiece, and action, to inspect/replace the tool.
@@
Update the hierarchical model if more data is available since the last iteration. Use the model’s predictions to calculate expected utilities at the current inspection point. is $P(T_{f}<T)\geq\alpha$? Replace the tool is $EU(action)\geq EU(inaction)$? Conduct inspection. Continue to next inspection point YesNoYesNo
@@
As mentioned previously, the first four tools are to be seen as historic data that the manufacturer has collected from previous tools. The hierarchical model can leverage this data to inform the population-level distributions. The active-learning procedure tool replacements are not calculated for these tools, only Tools 5-7. During the original experiment the measurements were taken at a time step that relates to 6.02 km sliding distance Wickramarachchi, (2019). Figure 10 shows the optimal replacement for Tools 5-7 based on the "gold standard" model.
@@ 4 Results
Table 2: A table containing the inspection values.
Tool replaced at (0.602 kms) Discrepancy (0.602 kms) Number of Inspections Inspection Type Tool 5 Tool 6 Tool 7 Tool 5 Tool 6 Tool 7 Full Information 10 10 11 0 0 0 0 Periodic Inspections 9 9 10 1 1 1 10 Risk-based Inspections 9 9 9 1 1 2 5
Table 3 compares the performance of the inspection approaches. The ’Cost of inspections’ column is the number of inspections multiplied by the cost of an inspection. The ’Cost of Wasted Tool Life’ for each inspection approach can be seen in the first column of Figure 3 and can be calculated using Equation 27.
@@
Table 3: A table containing the costs of each inspection method.
Cost of Wasted Tool life Cost of inspections Total Cost Inspection Type Tool 5 Tool 6 Tool 7 Tool 5 Tool 6 Tool 7 Periodic Inspections 0.025 0.025 0.028 0.15 0.15 0.15 0.573 Risk-based Inspections 0.025 0.025 0.056 0.15 0.05 0.05 0.361

### v2

@@ 3.3   Decision Process and Active Learning
A common quality-control criterion for the machining of engineering components is maintaining a surface roughness below some threshold level $S_{\text{crit}}$, beyond which the component is no longer fit for purpose. A high-quality surface finish can significantly improve the fatigue strength, corrosion resistance and creep life of machined parts [32]. If the surface finish is damaged via inadequate modelling or control of the machining process, the part may need to be discarded or re-machined; this has an associated utility, $C_{\text{workpiece}}$. Additionally, during the machining process an inspection can be conducted to gain access to a noisy observation of the surface roughness, with a utility $C_{\text{inspection}}$. Throughout the machining process the tool can be replaced, with a cost $C_{\text{tool}}$. Here, there is a decision, inaction, to allow the tool to continue machining with a risk to damage the workpiece, and action, to inspect/replace the tool.
@@
As mentioned previously, the first four tools are to be seen as historic data that the manufacturer has collected from previous tools. The hierarchical model can leverage this data to inform the population-level distributions. The active-learning procedure tool replacements are not calculated for these tools, only Tools 5-7. During the original experiment the measurements were taken at a time step that relates to 6.02 km sliding distance [35]. Figure 10 shows the optimal replacement for Tools 5-7 based on the "gold standard" model.
@@ 4   Results
Table 3 compares the performance of the inspection approaches. The ’Cost of inspections’ column is the number of inspections multiplied by the cost of an inspection. The ’Cost of Wasted Tool Life’ for each inspection approach can be seen in the first column of Figure LABEL:tab:costs and can be calculated using Equation 27.